If the demand function of a commodity is Q = 36 − 4P, where P and Q are price and quantity respectively, determine the price elasticity of demand when the price is R5. Indicate whether demand is elastic or inelastic at this price and provide justification for your answer. [1] εd = −1,25; because | − 1,25| = 1,25 > 1, demand is elastic [2] εd = −0,25; because | − 0,25| = −0,25 < 1, demand is elastic [3] εd = 0,25; because |0,25| = 0,25 < 1, demand is inelastic [4] εd = 1,25; because |1,25| = 1,25 > 1, demand is inelastic 


1
Expert's answer
2021-05-26T15:07:49-0400

Given that,

Demand Function is "Q = 36-4P" and

Price = 5


"Q = 36 - 4 \times 5"

  "\ \ \ = 36 - 20"

"=16"


"\dfrac{∆Q}{∆P}= -4"


We know that,

Price Elasticity Of demand"=\dfrac{ P}{Q}\times \dfrac{\delta Q}{\delta P}"

"=\dfrac{5}{16}x-4"

"=-1.25"


Hence the option A is correct

""\u03b5d = \u22121,25" ; because "| \u2212 1,25| = 1,25 > 1" , demand is elastic"